solve using the square root property : (x+9)^2=24
Answer by TimothyLamb(4379)

---
(x + 9)^2 = 24
(x + 9)(x + 9) = 24
xx + 18x + 81 = 24
xx + 18x + 57 = 0
---
the above quadratic equation is in standard form, with a=1, b=18, and c=57
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 18 57
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
answer:
x = -4.10102051
x = -13.8989795
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php